# Observing hidden neuronal states in experiments

### v2

@@ Abstract
Figure S5: (a) The representative voltage-clamp recordings of the responses to the slow ramps of the command potential. Top traces represent the whole-cell current, bottom trace represent the command potential protocol. Dark green lines correspond to the ascending ramp; light green lines correspond to the descending ramp. (b) Pairs of averaged ascending and descending I-V relationships, obtained in control conditions (left panel), with QX314 in the pipette solution (middle panel), and with both QX314 in the pipette solution and Cd2+ in the ACSF (right panel). Gray bars represent the standard error. The I-V relationships were normalized to the value of current at 0 mV. C. The comparison of I-V slope ratios in three experimental conditions (Kruskal-Wallis ANOVA, P$<0.001$, followed by Dunn’t post hoc test; - P$<0.05$, ∗∗ - P$<0.01$)
@@
We hypothesized that the mismatch between the ascending and descending I-V relationships can be attributed to the activity of the slow inactivating sodium and calcium ion channels. These slowly inactivating sodium and calcium channels are active at the ascending ramp, whereas high depolarization at the descending ramp inactivates them such that they are ineffective and thus do not contribute to the S-shaped I-V relationship. It has previously been reported that persistent sodium channels are present in deep entorhinal cortex neurons, and that current through these channels can be detected even when the rate of membrane depolarization is very low [44]. A prepulse to depolarized voltages inactivates persistent sodium channels of the neocortical neurons [40]. T-Type calcium channels are also known to produce the “window” currents due to the overlap of the activation and steady-state inactivation curves [45]. In order to check this hypothesis we used QX314 (6 mM) and Cd2+ (50 µM), which are the blockers of the voltage-gated sodium and calcium ion channels, respectively. To quantify the effects of the applied substances on the mismatch between the ascending and descending I-V relationships we introduced the I-V slope ratio (Fig. S5 (c)). It was calculated as the ratio of the minimal slope of the ascending I-V relationship to the slope of the descending I-V relationship at -35 mV. In control conditions, the I-V slope ratio was negative due to a strong mismatch between the I-V relationships (Fig. S5 (b,c), CTRL). When the pipette solution containing QX314 was used, the observed mismatch between the I-V relationships was much less pronounced (Fig. S5 (b), middle panel), resulting in the significantly higher I-V slope ratio. With QX314 in the pipette solution and Cd2+ in the ACSF, both the ascending and descending I-V relationships were monotonically increasing functions (Fig. S5 (b), right panel), resulting in positive values of the I-V slope ratios. The results obtained suggest that the negative slope of the ascending I-V relationship is mainly mediated by persistent sodium channels and to a lesser extent by voltage-gated calcium channels.
@@
In the model the channel properties depend sensitively on the maximal conductances (e.g., $g_{\mathrm{Ca}}m_{\infty}$ for calcium in eq. (8) of the main manuscript). A small change in these parameters in eq. (8) removes the two fold points from the steady-state curve (in a so-called cusp bifurcation), resulting in a change from type-I to type-II behavior.
@@
where (.)(.)( . ) means element-wise multiplication (so, $v.w=\operatorname{diag}(v)w$), and $\|v\|_{1}:=|v_{1}|+\ldots+|v_{n_{\mathrm{c}}}|$ is the $1$-norm of a vector $v$. The condition $m_{\mathrm{c}}<1$ is also sufficient for linear stability of $(V_{\mathrm{eq}},x_{\mathrm{eq},(\cdot)})$: the Jacobian of (S2) in $(V_{\mathrm{eq}},x_{\mathrm{eq},(\cdot)})$ is
@@
(note that the matrix entries are in row order $1\times 1$, $1\times{n_{\mathrm{c}}}$, ${n_{\mathrm{c}}}\times 1$ and ${n_{\mathrm{c}}}\times{n_{\mathrm{c}}}$, and that the term $\operatorname{diag}(r_{\mathrm{eq}}^{\prime})x_{\infty}(V_{\mathrm{eq}})$ cancels in the bottom left entry). By rescaling the $x_{j}$ to $x_{j}/(x_{\infty,j}^{\prime}(1+\delta)$ with an arbitrary $\delta>0$ we can make the lower ${n_{\mathrm{c}}}$ rows diagonally dominant. This gives then in the upper right entry $-(I_{x}.x_{\infty}^{\prime})^{\mathsf{T}}(1+\delta)/C$:
@@
where $I_{\mathrm{vc}}(V_{\mathrm{h}})$ is the measured $(I,V)$-curve during voltage clamp runs (VC) and $\tau_{1/2}$ is the time of decay to half of the transients during large fluctuations of measured $I_{\mathrm{vc}}$ during VC runs (see Section S3 and Fig. S4). The estimates (S5) are themselves accurate to first order in $\varepsilon$, such that their bias is only a second-order effect in (S4) if $\alpha>0$. Pretending $I_{\mathrm{vc}}(t)$ was governed by a scalar first-order ordinary differential equation $\dot{I}_{\mathrm{vc}}=f_{IV}(I_{\mathrm{vc}},V_{\mathrm{h}})$, bias estimate (S4) would follow from implicit differentiation with respect to time and $I_{\mathrm{vc}}$ , where $\alpha=-\partial_{1}f_{IV}(I_{\mathrm{eq}}(V_{\mathrm{h}}),V_{\mathrm{h}})>0$.
@@
Examples of such models are equations (1), (2) in the main text; see also equation (8) for the concrete Morris-Lecar model used for Fig. 2 in the main text and Fig. S4. More precisely, let us collect the entire vector of conductance-based model variables into $y=(V,x_{j})_{j=1,\ldots,n_{\mathrm{c}}}$, such that the conductance-based model (1), (2) for voltage clamp with a slow ramp of hold voltage $V_{\mathrm{h}}$ from the main text is of the general form
@@
The first condition for validity of (S4) is that among the eigenvalues of the matrix $J(V_{\mathrm{h}}):=\partial_{1}f(y_{\mathrm{eq}}(V_{\mathrm{h}}),V_{\mathrm{h}})$ (which all have negative part), there is one dominant eigenvalue $-\alpha(V_{\mathrm{h}})$ of $J(V_{\mathrm{h}})$ with largest real part. This dominance is expressed as a spectral gap condition for $J(V_{\mathrm{h}})$, which means that all other eigenvalues must have real part less than some bound $-\beta(V_{\mathrm{h}})<-\alpha(V_{\mathrm{h}})<0$ with the condition
$\beta^{-1}(V_{\mathrm{h}})$ $\ll\alpha^{-1}(V_{\mathrm{h}})\sim\mbox{\ order 1.}$. (S7)
@@
$I_{\mathrm{vc}}(y)\approx I_{\mathrm{vc}}(y_{\mathrm{eq}}(V_{\mathrm{h}}))+\partial I_{\mathrm{vc}}(y_{\mathrm{eq}}(V_{\mathrm{h}}))[y-y_{\mathrm{eq}}(V_{\mathrm{h}})]\mbox{,\quad where $\partial I_{\mathrm{vc}}(y_{\mathrm{eq}}(V_{\mathrm{h}}))=:c(V_{\mathrm{h}})^{\mathsf{T}}\in\mathbb{R}^{1\times(n_{\mathrm{c}}+1)}$ is a row vector.}$.
@@
The second condition for validity of (S4) is that the right eigenvector $y_{\mathrm{ev}}(V_{\mathrm{h}})$ for the dominant eigenvalue $-\alpha(V_{\mathrm{h}})$ has a non-zero (more precisely, non-small) projection in variations of this measured output:
@@
If the dominant direction $y_{\mathrm{ev}}$ is observable according to (S8), then we may normalize the eigenvector $y_{\mathrm{ev}}$ such that $c(V_{\mathrm{h}})^{\mathsf{T}}y_{\mathrm{ev}}(V_{\mathrm{h}})=1$ for all $V_{\mathrm{h}}$. We denote the left eigenvector of $J(V_{\mathrm{h}})$ for the dominant eigenvalue $-\alpha(V_{\mathrm{h}})$ by $y_{\mathrm{ad}}(V_{\mathrm{h}})$, normalized such that $y_{\mathrm{ad}}(V_{\mathrm{h}})^{\mathsf{T}}y_{\mathrm{ev}}(V_{\mathrm{h}})=1$. If $J(V_{\mathrm{h}})$ satisfies the spectral gap condition (S7), its inverse is $J(V_{\mathrm{h}})^{-1}\approx-\alpha(V_{\mathrm{h}})y_{\mathrm{ev}}(V_{\mathrm{h}})y_{\mathrm{ev}}(V_{\mathrm{h}})^{\mathsf{T}}$ (ignoring small terms of order $\beta^{-1}$), thus, it is close to the rank-$1$ matrix $-\alpha y_{\mathrm{ev}}y_{\mathrm{ad}}^{\mathsf{T}}$. As $y_{\mathrm{eq}}(V_{\mathrm{h}})$ is the steady-state curve of (S6) (so, $0=f(y_{\mathrm{eq}}(V_{\mathrm{h}}),V_{\mathrm{h}})$), we have (abbreviating $\partial_{2}f_{\mathrm{eq}}(V_{\mathrm{h}}):=\partial_{2}f(y_{\mathrm{eq}}(V_{\mathrm{h}}),V_{\mathrm{h}})$ and dropping argument $V_{\mathrm{h}}$ from $\alpha$, $c$, $y_{\mathrm{ev}}$, $y_{\mathrm{ad}}$, $J$, and $\partial_{2}f_{\mathrm{eq}}$)
[…]
@@
Finally, we observe that the measured current $I_{\mathrm{vc}}(t)$ when $V_{\mathrm{h}}$ changes dynamically is not $-\sum_{j}I_{j}(x_{j},V)$ but (as measured at the electrode) equal to $g_{\mathrm{c}}(V_{\mathrm{h}}-V)$. However, $I_{\mathrm{vc}}(t)=g_{\mathrm{c}}(V_{\mathrm{h}}(t)-V(t))=\sum_{j}I_{j}(x_{j},V)+C\dot{V}(t)$, where $\dot{V}$ is a component of $\dot{y}\approx f(y,V_{\mathrm{h}})\approx J[y(t)-y_{\mathrm{eq}}(V_{\mathrm{h}}(t))]$, which equals $y_{\mathrm{eq}}^{\prime}(V_{\mathrm{h}})\dot{V}_{\mathrm{h}}$ and is of order $\varepsilon$ after transients have decayed. Thus, the correction due to this difference in (S9) is only another $O(\varepsilon)$ term in $I_{\mathrm{eq}}^{\prime}(V_{\mathrm{h}})$, making it a second-order effect overall.

### v3

@@ Acknowledgments
Figure E in S1 Text: A: The representative voltage-clamp recordings of the responses to the slow ramps of the command potential. Top traces represent the whole-cell current, bottom trace represent the command potential protocol. Dark green lines correspond to the ascending ramp; light green lines correspond to the descending ramp. B: Pairs of averaged ascending and descending I-V relationships, obtained in control conditions (left panel), with QX314 in the pipette solution (middle panel), and with both QX314 in the pipette solution and Cd2+ in the ACSF (right panel). Gray bars represent the standard error. The I-V relationships were normalized to the value of current at 0 mV. C: The comparison of I-V slope ratios in three experimental conditions (Kruskal-Wallis ANOVA, P<0.001<0.001, followed by Dunn’t post hoc test; - P<0.05<0.05, ∗∗ - P<0.01<0.01)
@@
We hypothesized that the mismatch between the ascending and descending I-V relationships can be attributed to the activity of the slowly inactivating sodium and calcium ion channels. These slowly inactivating sodium and calcium channels are active at the ascending ramp, whereas high depolarization at the descending ramp inactivates them such that they are ineffective and thus do not contribute to the S-shaped I-V relationship. It has previously been reported that persistent sodium channels are present in deep entorhinal cortex neurons, and that current through these channels can be detected even when the rate of membrane depolarization is very low [agrawal2001]. A prepulse to depolarized voltages inactivates persistent sodium channels of the neocortical neurons [fleidervish1996]. T-Type calcium channels are also known to produce the “window” currents due to the overlap of the activation and steady-state inactivation curves [amarillo2014]. In order to check this hypothesis we used QX314 (6 mM) and Cd2+ (50 µM), which are the blockers of the voltage-gated sodium and calcium ion channels, respectively. To quantify the effects of the applied substances on the mismatch between the ascending and descending I-V relationships we introduced the I-V slope ratio (Fig EC in S1 Text). It was calculated as the ratio of the minimal slope of the ascending I-V relationship to the slope of the descending I-V relationship at -35 mV. In control conditions, the I-V slope ratio was negative due to a strong mismatch between the I-V relationships (Fig EB and EC in S1 Text, CTRL). When the pipette solution containing QX314 was used, the observed mismatch between the I-V relationships was much less pronounced (Fig EB in S1 Text, middle panel), resulting in the significantly higher I-V slope ratio. With QX314 in the pipette solution and Cd2+ in the ACSF, both the ascending and descending I-V relationships were monotonically increasing functions (Fig EB in S1 Text, right panel), resulting in positive values of the I-V slope ratios. The results obtained suggest that the negative slope of the ascending I-V relationship is mainly mediated by persistent sodium channels and to a lesser extent by voltage-gated calcium channels. To avoid inactivation of these channels by high depolarization all voltage-clamp runs in the main manuscript increase the hold voltage $V_{\mathrm{h}}$, and all current-clamp runs increase the hold current $I_{\mathrm{h}}$.
@@
In the model the channel properties depend sensitively on the maximal conductances (e.g., $g_{\mathrm{Ca}}m_{\infty}$ for calcium in eq (8) of the main manuscript). A small change in these parameters in eq. (8) removes the two fold points from the steady-state curve (in a so-called cusp bifurcation), resulting in a change from class-I to class-II behavior.
@@
where $(.)$ means element-wise multiplication (so, $v.w=\operatorname{diag}(v)w$), and $\|v\|_{1}:=|v_{1}|+\ldots+|v_{n_{\mathrm{c}}}|$ is the $1$-norm of a vector $v$. The condition $m_{\mathrm{c}}<1$ is also sufficient for linear stability of $(V_{\mathrm{eq}},x_{\mathrm{eq},(\cdot)})$: the Jacobian of (S2) in $(V_{\mathrm{eq}},x_{\mathrm{eq},(\cdot)})$ is
@@
(note that the matrix entries are in row order $1\times 1$, $1\times{n_{\mathrm{c}}}$, ${n_{\mathrm{c}}}\times 1$ and ${n_{\mathrm{c}}}\times{n_{\mathrm{c}}}$, and that the term $\operatorname{diag}(r_{\mathrm{eq}}^{\prime})x_{\infty}(V_{\mathrm{eq}})$ cancels in the bottom left entry). By rescaling the $x_{j}$ to $x_{j}/(x_{\infty,j}^{\prime}(1+\delta))$ with an arbitrary $\delta>0$ we can make the lower $n_{\mathrm{c}}$} rows diagonally dominant. This gives then in the upper right entry $(I_{x}.x_{\infty}^{\prime})^{\mathsf{T}}(1+\delta)/C$:
@@
where $I_{\mathrm{vc}}(V_{\mathrm{h}})$ is the measured $(I,V)$-curve during voltage clamp runs (VC) and $\tau_{1/2}$ is the time of decay to half of the transients during large fluctuations of measured $I_{\mathrm{vc}}$ during VC runs (see Section S3 and Fig D in S1 Text). The estimates (S5) are themselves accurate to first order in $\varepsilon$, such that their bias is only a second-order effect in (S4) if $\alpha>0$. Pretending $I_{\mathrm{vc}}(t)$ was governed by a scalar first-order ordinary differential equation $\dot{I}_{\mathrm{vc}}=f_{IV}(I_{\mathrm{vc}},V_{\mathrm{h}})$, bias estimate (S4) would follow from implicit differentiation with respect to time and $I_{\mathrm{vc}}$ , where $\alpha=-\partial_{1}f_{IV}(I_{\mathrm{eq}}(V_{\mathrm{h}}),V_{\mathrm{h}})>0$.
@@
Examples of such models are equations (1), (2) in the main text; see also the Morris-Lecar model used for Fig 3 of the main manuscript and Fig D in S1 Text, given in Materials and Methods in the main manuscript. More precisely, let us collect the entire vector of conductance-based model variables into $y=(V,x_{j})_{j=1,\ldots,n_{\mathrm{c}}}$, such that the conductance-based model (1), (2) for voltage clamp with a slow ramp of hold voltage $V_{\mathrm{h}}$ from the main text is of the general form
@@
The first condition for validity of (S4) is that among the eigenvalues of the matrix $J(V_{\mathrm{h}}):=\partial_{1}f(y_{\mathrm{eq}}(V_{\mathrm{h}}),V_{\mathrm{h}})$ (which all have negative part), there is one dominant eigenvalue $\alpha(V_{\mathrm{h}})$ of $J(V_{\mathrm{h}})$ with largest real part. This dominance is expressed as a spectral gap condition for $J(V_{\mathrm{h}})$, which means that all other eigenvalues must have real part less than some bound $\beta(V_{\mathrm{h}})<-\alpha(V_{\mathrm{h}})<0$ with the condition
$\beta^{-1}(V_{\mathrm{h}})$ $\ll\alpha^{-1}(V_{\mathrm{h}})\sim\mbox{\ order 1.}$ (S7)
@@
$I_{\mathrm{vc}}(y)\approx I_{\mathrm{vc}}(y_{\mathrm{eq}}(V_{\mathrm{h}}))+\partial I_{\mathrm{vc}}(y_{\mathrm{eq}}(V_{\mathrm{h}}))[y-y_{\mathrm{eq}}(V_{\mathrm{h}})]\mbox{,\quad where $\partial I_{\mathrm{vc}}(y_{\mathrm{eq}}(V_{\mathrm{h}}))=:c(V_{\mathrm{h}})^{\mathsf{T}}\in\mathbb{R}^{1\times(n_{\mathrm{c}}+1)}$ is a row vector.}$
@@
The second condition for validity of (S4) is that the right eigenvector $y_{\mathrm{ev}}(V_{\mathrm{h}})$ for the dominant eigenvalue $\alpha(V_{\mathrm{h}})$ has a non-zero (more precisely, non-small) projection in variations of this measured output:
@@
If the dominant direction $y_{\mathrm{ev}}$ is observable according to (S8), then we may normalize the eigenvector $y_{\mathrm{ev}}$ such that $c(V_{\mathrm{h}})^{\mathsf{T}}y_{\mathrm{ev}}(V_{\mathrm{h}})=1$ for all $V_{\mathrm{h}}$. We denote the left eigenvector of $J(V_{\mathrm{h}})$ for the dominant eigenvalue $\alpha(V_{\mathrm{h}})$ by $y_{\mathrm{ad}}(V_{\mathrm{h}})$, normalized such that $y_{\mathrm{ad}}(V_{\mathrm{h}})^{\mathsf{T}}y_{\mathrm{ev}}(V_{\mathrm{h}})=1$. If $J(V_{\mathrm{h}})$ satisfies the spectral gap condition (S7), its inverse is $J(V_{\mathrm{h}})^{-1}\approx-\alpha(V_{\mathrm{h}})y_{\mathrm{ev}}(V_{\mathrm{h}})y_{\mathrm{ev}}(V_{\mathrm{h}})^{\mathsf{T}}$ (ignoring small terms of order $\beta^{-1}$), thus, it is close to the rank-$1$ matrix $\alpha y_{\mathrm{ev}}y_{\mathrm{ad}}^{\mathsf{T}}$. As $y_{\mathrm{eq}}(V_{\mathrm{h}})$ is the steady-state curve of (S6) (so, $0=f(y_{\mathrm{eq}}(V_{\mathrm{h}}),V_{\mathrm{h}})$), we have (abbreviating $\partial_{2}f_{\mathrm{eq}}(V_{\mathrm{h}}):=\partial_{2}f(y_{\mathrm{eq}}(V_{\mathrm{h}}),V_{\mathrm{h}})$ and dropping argument $V_{\mathrm{h}}$ from $\alpha$, $c$, $y_{\mathrm{ev}}$, $y_{\mathrm{ad}}$, $J$, and $\partial_{2}f_{\mathrm{eq}}$)
$y_{\mathrm{eq}}^{\prime}(V_{\mathrm{h}})$ $=-J^{-1}\partial_{2}f_{\mathrm{eq}}\approx-\alpha y_{\mathrm{ev}}y_{\mathrm{ad}}^{\mathsf{T}}\partial_{2}f_{\mathrm{eq}}\mbox{, \quad and, hence,}$
@@
Finally, we observe that the measured current $I_{\mathrm{vc}}(t)$ when $V_{\mathrm{h}}$ changes dynamically is not $\sum_{j}I_{j}(x_{j},V)$ but (as measured at the electrode) equal to $g_{\mathrm{c}}(V_{\mathrm{h}}-V)$. However, $I_{\mathrm{vc}}(t)=g_{\mathrm{c}}(V_{\mathrm{h}}(t)-V(t))=\sum_{j}I_{j}(x_{j},V)+C\dot{V}(t)$, where $\dot{V}$ is a component of $\dot{y}\approx f(y,V_{\mathrm{h}})\approx J[y(t)-y_{\mathrm{eq}}(V_{\mathrm{h}}(t))]$, which equals $y_{\mathrm{eq}}^{\prime}(V_{\mathrm{h}})\dot{V}_{\mathrm{h}}$ and is of order $\varepsilon$ after transients have decayed. Thus, the correction due to this difference in (S9) is only another $O(\varepsilon)$ term in $I_{\mathrm{eq}}^{\prime}(V_{\mathrm{h}})$, making it a second-order effect overall.
S7 Numerical bifurcation diagrams for in-silico models
Figure F in S1 Text: Numerical bifurcation diagram overlaid for Fig 3A from the main manuscript, showing a class-I Morris-Lecar neuron model [morris1981] as example of excitatory cell. Black curves show $\max_{t}V(t)$ and $\min_{t}V(t)$ for periodic firing oscillations at each fixed $I_{\mathrm{h}}$ (solid for dynamically stable, dashed for dynamically unstable). Equations and parameter values are given in Material and Methods of the main manuscript. The diagrams on the right show the dependence of the period of the periodic orbits on the parameter and a typical periodic time profile for $I_{\mathrm{h}}=100$ pA.
Figure G in S1 Text: Numerical bifurcation diagram overlaid for Fig 3B from the main manuscript, showing a Wang-Buzsaki neuron model [wang1996] as example of inhibitory interneuron. Black curves show $\max_{t}V(t)$ and $\min_{t}V(t)$ for periodic firing oscillations at each fixed $I_{\mathrm{h}}$ (solid for dynamically stable, dashed for dynamically unstable). Equations and parameter values are given in Material and Methods of the main manuscript. The diagrams on the right show the dependence of the period of the periodic orbits on the parameter and a typical periodic time profile for $I_{\mathrm{h}}=10$ pA.
Figure H in S1 Text: Numerical bifurcation diagram overlaid for Fig 4 from the main manuscript, showing a class-II Morris-Lecar neuron model [morris1981] as example of class-II neuron. Black curves show $\max_{t}V(t)$ and $\min_{t}V(t)$ for periodic firing oscillations at each fixed $I_{\mathrm{h}}$ (solid for dynamically stable, dashed for dynamically unstable). Equations and parameter values are given in Material and Methods of the main manuscript. The diagrams on the right show the dependence of the period of the periodic orbits on the parameter and a typical periodic time profile for $I_{\mathrm{h}}=250$ pA.
The diagrams in Figs F, G and H in S1 Text contain copies of the graphs of VC and CC protocols applied to the Morris-Lecar model [morris1981] and the Wang-Buzaki model [wang1996], as shown in Figs 3 and 4 of the main manuscript. All equations and parameter values are given in Materials and Methods of the main manuscript. Figs F, G and H in S1 Text also contain overlaid numerical bifurcation diagrams obtained using coco [DS13]. The code reproducing the figures is available at https://github.com/serafimrodrigues/Amakhin-etal-2025. The repository includes a snapshot of coco (revision 3328) as downloaded from https://sourceforge.net/projects/cocotools.
Near Hopf bifurcations at $I_{\mathrm{db}}$ in Fig F in S1 Text and at $I_{\mathrm{h}}\approx 135$ pA in Fig H in S1 Text we observe a small region of bistability. This region shrinks for stronger time scale separation (parameter $\phi$ in the Morris-Lecar model).
S8 Zoom near $I_{\mathrm{db}}$ for CC protocol of cell 5
Figure I in S1 Text: Zoom of CC run for cell 5 near the estimate for $I_{\mathrm{db}}\approx 365$ pA. Parts of the recordings where smoothed time profiles of $(V_{\mathrm{cc}}^{\prime}(t),V_{\mathrm{cc}}^{\prime\prime}(t))$ have a norm smaller than a threshold are highlighted by a thicker line and labelled as slow part of the firing oscillation or near equilibrium.
Fig I in S1 Text shows a zoom of the CC run data for cell 5 of Fig 1B of the main manuscript.
Highlighting of slow part of voltage profile in CC run
The criterion for labelling part of the time series $V_{\mathrm{cc}}$ of a CC run as slow is if the time profile $(V_{\mathrm{cc}}^{\prime},V_{\mathrm{cc}}^{\prime\prime})$ has a norm less than some threshold $\theta_{\mathrm{th}}$ at time $t_{i}$ for sampling times $t_{i}$. More precisely, let $(t_{i},V_{i})$ be the voltages $V_{i}$ sampled at times $t_{i}$ for $i=1,\ldots,n_{\mathrm{s}}$. We define the operations
$t^{(k)}$ = $=\mbox{\ sampling times $t$, refined by $k$ uniformly spaced }$
 intermediate times between $t_{i-1}$ and $t_{i}$ for each $i=2,\ldots,n_{\mathrm{s}}$,
$V^{(k)}$ = $=\mbox{\ linear interpolation of $(t_{i},V_{i})$ at $t^{(k)}$,}$
$\operatorname{sm}_{w}x$ $=\mathtt{smoothdata(}x\mathtt{,"gaussian",}w\mathtt{,"SamplePoints",}t^{(k)}\mathtt{)}$
$\operatorname{sp}x$ $=\mathtt{interp1(}t^{(k)},x\mathtt{,"spline","pp")}$
$\operatorname{ev}f$ $=f(t^{(k)})$
(with the Matlab commands smoothdata for smoothing, interp1 for cubic spline interpolation), where we note that $\operatorname{sp}$ is applied to a profile (vector) sampled at $t^{(k)}$, resulting in a differentiable function, while $\operatorname{ev}$ applies a function on the sample points $t^{(k)}$ resulting in a vector sampled on $t^{(k)}$. With these operations we construct a smoothed differentiation for a profile $x$ on $t^{(k)}$ as
$x^{\prime}$ $=\operatorname{ev}\circ\frac{\mathrm{d}}{\mathrm{d}t}\circ\operatorname{sp}\circ\operatorname{sm}_{w}^{2}x.$
With this definition $V_{\mathrm{cc}}$ and $I_{\mathrm{h}}$ at all times $t^{(k)}_{i}$ for which
$\left({V^{(k)}}^{\prime\prime}\right)^{2}\Delta_{t}^{2}+\left({V^{(k)}}^{\prime}\right)^{2}\Delta_{t}$ $<\theta_{\mathrm{th}}^{2}$
as slow (thick line) in Figs I in S1 Text, 1B and 2 of the main manuscript. Parameter values were $k=5$, $w=0.01$ s, $\theta_{\mathrm{th}}=5$ for Figs 1B, 2, A and I in S1 Text. The refining by $k$ additional points in for the sampling $t_{i}$ enables the application of the operation $\operatorname{sm}_{w}$ with a window size close to the sampling time.